Find the integral curves of the differential equation:


(D³-D'³) z=x³y³

1
Expert's answer
2022-01-25T14:51:02-0500

The given equation can be written as

"(D^3 \u2013 D'^3)z = x^3y^3" --------(1)

Corresponding auxiliary equation is

"m^3 \u2013 1 = 0" , i.e "(m \u2013 1)(m^2 + m + 1) = 0"

or

"(m \u2013 1)(m \u2013 w)(m \u2013 w^2) = 0," for w to be the cube root of unity----(2)

"\displaystyle C.F. = φ_{1}(y + x) + φ_2(y + wx) + φ_3(y + w^2x)" -----(3)

Now

"\displaystyle p.I=\frac{1}{D^3-D^{\prime3}}x^3y^3=\frac{1}{D^3\left[1-(\frac{D\prime}{D})^3\right]}x^3y^3"


"=\frac{1}{D^3}\left[1-(\frac{D\prime}{D})^3\right]^{-1}x^3y^3=\frac{1}{D^3}\left[1+\frac{D^\prime3}{D}+...\right]x^3y^3"


"=\frac{1}{D^3}x^3y^3+\frac{D^\prime3}{D^6}x^3y^3"


(on taking "\frac{D^\prime}{D}"  to the powers to which y is appearing in

"F(x, y) = x^ny^m" i.e neglecting terms containing powers more than 3 in "D^\prime"

"\displaystyle =\frac{x^6y^3}{4\cdot5\cdot6}+\frac{6x^9}{4\cdot5\cdot6\cdot7\cdot8\cdot9}"


"P.I\displaystyle =\frac{x^6y^3}{ 120}+\frac{x^9}{ 10080}"

Hence complete solution 

"\displaystyle z=φ_{1}(y+x)+φ_{2}(y+wx)+φ_{3}(y+w^2x)+\frac{x^6y^3}{ 120}+\frac{x^9}{ 10080}"